Rogelio and jin are looking up at the top of a tower that 36.6m tall. If jin is looking at an angle of elevation of 24 degrees, and rogelio is on the opposite side of the tower looking up at an angle of 17 degree, how far apart are rogelio and jin to the nearest meter? Assume their eyes are both 1.8m off the ground.
Answer by dfrazzetto(283)

They each form a right triangle with side 'a' being the height of the tower minus their height (to the eye), and side 'b' being their distance from the tower
a = 36.6 - 1.8m = 34.8m
tan(angle) = a/b
Jin:
tan(24) = 34.8/b
b = 34.8 / tan(24) = 78.16
Rogelio:
tan(17) = 34.8/b
b = 34.8/ tan(17) = 113.83
Total distance apart = b(jin) + b(rogelio) = 78.16 + 113.83 = 191.99 = 192 m